### v1

@@ Abstract
Recent advancements in machine learning have led to novel imaging systems and algorithms that address ill-posed problems. Assessing their trustworthiness and understanding how to deploy them safely at test time remains an important and open problem. We propose a method that leverages conformal prediction to retrieve upper/lower bounds and statistical inliers/outliers of reconstructions based on the prediction intervals of downstream metrics. We apply our method to sparse-view CT for downstream radiotherapy planning and show 1) that metric-guided bounds have valid coverage for downstream metrics while conventional pixel-wise bounds do not and 2) anatomical differences of upper/lower bounds between metric-guided and pixel-wise methods. Our work paves the way for more meaningful reconstruction bounds. Code available at https://github.com/matthewyccheung/conformal-metric.
Keywords:
Uncertainty Conformal Prediction Inverse problems
@@ 1 Introduction
Recent advancements in machine learning have led to novel imaging systems and algorithms that address ill-posed problems. Traditionally, image reconstruction evaluation relies on common image quality metrics such as PSNR, SSIM, FID, and Dice scores of segmentations on reconstructed images. While these metrics provide a heuristic to gauge the overall model uncertainty in reconstruction during evaluation, they do not provide uncertainties and guarantees at test time, and do not link reconstruction quality to uncertainties in downstream applications.
@@
Conformal prediction (CP) provides distribution-free, valid, and calibrated prediction intervals at test time [32, 10, 29, 2]. The idea is to use residuals from a calibration dataset to infer uncertainty in future test datasets [32, 10, 29, 2]. For regression tasks, this uncertainty is given as a prediction interval [32, 10, 29, 2]. The application of CP to image reconstruction has been relatively limited. This is a difficult problem because quantiles in higher dimensional data are M-quantiles, meaning they have infinite solutions and only have unique solutions when a direction is specified [6, 7]. How do we pick such a direction? The conventional (pixel-wise) method is to pick the direction where all pixels are independent [12, 16, 9]. The upper and lower bounds of the estimated image can be calibrated based on pixel intensity [3, 14]. While these pixel-wise prediction intervals are easy to interpret, they do not consider spatial correlations and may lead to large interval lengths [4]. The upper and lower bounds can also be calibrated in the direction of principal components [4]. While using principal components considers spatial correlations, it does not capture meaningful and practical uncertainty for downstream processes and is prohibitively costly to compute for large images. Furthermore, both methods provide upper and lower bounds not sampled from the learned manifold, yielding implausible images. A reasonable answer is to calibrate the upper and lower bounds in the direction of semantic features [27]. However, this method requires training a generative model with disentangled latent spaces.
We argue that bounds should be computed in the direction of downstream metrics for more reliable downstream performance. We propose “Metric-guided Image Reconstruction Bounds” that leverages CP to form valid prediction intervals of reconstructions in the direction of downstream metrics and retrieve reconstructions 1) closest to the upper and lower bounds, 2) contained in the bounds (statistical inliers) and 3) outside the bounds (statistical outliers). Our method takes spatial correlations into account and produces plausible reconstructions from the learned manifold. We show that our method provides valid coverage for downstream tasks while the conventional pixel-wise method does not and the upper/lower bounds between methods are anatomically different.
We demonstrate our method on sparse-view computed tomography (sv-CT) and downstream radiotherapy planning. Reconstruction is highly accurate for CT machines and uses sophisticated detectors and algorithms to obtain sub-millimeter spatial resolution. CT downtime significantly impacts the availability of radiotherapy planning in low-resource settings [11]. A low-cost device with cone-beam geometry could be manufactured and used to increase access to radiotherapy planning and other therapeutic use cases. Individualized radiotherapy plans are made from CT scans and specify localized doses to a target treatment volume (i.e. breast, prostate). We use downstream clinical application metrics from radiotherapy planning to retrieve reconstructions.
@@ 2 Method
We consider a 3-D reconstruction setting for a downstream application with a chosen downstream metric. [^1] While we concentrate on imaging, our method can be applied to any multidimensional setting. The measurement and reconstruction algorithms are assumed to be probabilistic. We follow the split conformal prediction procedure [26, 32, 2, 29] by using $n_{p}$ patients for the calibration dataset and 1 test patient $n_{p}+1$ with unknown ground truth volume and metric as the test dataset. For each patient $i$ in the calibration dataset, we reconstruct a set of volumes $\hat{V}^{i}=\{\hat{V}_{j}^{i}\}_{j=1}^{n_{r}}$ of size $n_{r}$. Each patient’s reconstructed volumes are used to attain a set of estimated metrics $\hat{Y}^{i}=\{\hat{Y}_{j}^{i}\}_{j=1}^{n_{r}}$. Each patient’s ground truth volume is used to attain a ground truth metric $Y^{i}$. For the test patient, we reconstruct a set of volumes $\hat{V}^{n_{p}+1}=\{\hat{V}_{j}^{n_{p}+1}\}_{j=1}^{n_{r}}$ and estimate metrics $\hat{Y}^{n_{p}+1}=\{\hat{Y}_{j}^{n_{p}+1}\}_{j=1}^{n_{r}}$. [^2] We do not have ground truth volume $V^{n_{p}+1}$ and metric $Y^{n_{p}+1}$ at test time. Assuming $(\hat{Y}^{i},Y^{i})$ for $i=1,...,n_{p}+1$ are exchangeable, we leverage Conformalized Quantile Regression (CQR) [26] to find the prediction interval $C(\hat{Y}^{n_{p}+1})$ satisfying the conformal coverage guarantee [33]:
$\mathbb{P}[Y^{n_{p}+1}\in C(\hat{Y}^{n_{p}+1})]\geq 1-\alpha$ (1)
where $\alpha$ is a user-specified miscoverage rate. We attain $C(\hat{Y}^{n_{p}+1})$ by adjusting the upper and lower bounds of $\hat{Y}^{n_{p}+1}$ with an offset $q$ that is computed from the calibration dataset to satisfy (1):
$C(\hat{Y}^{n_{p}+1})=[Q_{\alpha/2}(\hat{Y}^{n_{p}+1})-q,Q_{1-\alpha/2}(\hat{Y}^{n_{p}+1})+q]$ (2)
where $Q_{\alpha}(.)$ is the function that estimates the $\alpha$th quantile. [^3] While we use symmetric adjustments in this work, asymmetric bounds can also be used. See [26] for more details. Finally, we retrieve the volumes 1) closest to the upper and lower bounds of the prediction intervals $[\hat{V}_{LB}^{n_{p}+1},\hat{V}_{UB}^{n_{p}+1}]$ based on the $L_{1}$ norm, 2) contained within the prediction intervals (inliers), and 3) outside the prediction intervals (outliers). We provide an overview in Fig. 1 and pseudo-code in Algorithm 1.
Figure 1: Overview of our approach. Assume probabilistic measurement and reconstruction processes, $n_{p}$ patients for calibration, and 1 patient for testing. For test patient $n_{p}+1$ with unknown ground truth reconstruction and metric, 1) acquire measurements, 2) attain a set of reconstructions, 3) extract downstream metrics, 4) adjust upper and lower bounds of metric based on a calibration procedure, and 5) retrieve reconstructions with metrics closest to the calibrated upper and lower bounds, contained within bounds (statistical inliers), and outside of bounds (statistical outliers).
Algorithm 1 Metric-guided Image Reconstruction Bounds
[…]
@@
$q=Q_{\frac{\lceil(n_{p}+1)(1-\alpha)\rceil}{n_{p}}}(scores)$
[…]
@@
[…]
@@
[…]
for $j=1:n_{r}$ do
     if $\hat{Y}_{j}^{n_{p}+1}\in[LB(\hat{Y}^{n_{p}+1}),UB(\hat{Y}^{n_{p}+1})]$ then
         Add $\hat{Y}_{j}^{n_{p}+1}$ to inliers
     else
         Add $\hat{Y}_{j}^{n_{p}+1}$ to outliers
     end if
end for
Similar to prior work [4], we use sample quantiles instead of regression quantiles. Our method can be interpreted as a discrete version of CQR that finds marginal prediction intervals for downstream metrics given a patient. Our method is different to prior work in pixel-wise [3] that produce prediction intervals per reconstruction. Instead, we provide prediction sets directly from a set of patient reconstructions where each patient has different reconstruction volume sizes. We compare our method with conventional pixel-wise bounds.
@@ 3 Experiments
Dataset: We use a de-identified CT dataset of 20 patients retrospectively treated with radiotherapy at The University of Texas MD Anderson Cancer Center. All CT images were of patients who had received surgical mastectomy to the right side of the body, and radiotherapy to the post-mastectomy chest wall and/or axillary lymph nodes. This research was conducted using an approved institutional review board protocol. Each ground truth CT is of size ($\times 512\times$ 512 × Number of slices). For each patient, we generate 10 digitally reconstructed radiographs (DRR) from the ground truth CT scan using the TIGRE toolbox [5]. The DRRs simulate image acquisition from a cone-beam geometry. We simulate physical randomness (beam angle variability and sensor noise) by generating DRRs with 3% noise and 50 random projections between 0 and 360 degrees. The number of projections was increased from 2 to 50 until organ boundaries were perceptually discernible in the reconstruction by the RPA. Because this work aims to showcase the feasibility of CP for image reconstruction, we assume that such a low-cost sv-CT device will be created in future work that gives acceptable reconstruction image quality. We use a self-supervised model, Neural Attenuation Fields (NAF), for reconstruction [35]. Each reconstruction is uncropped and contains the full scan. We use the default parameter setting in NAF [35] and introduce computational randomness through random initializations of NAF [30, 19]. Ultimately, we construct a dataset of 20 patients with 10 reconstructions each. To construct the conventional pixel-wise upper and lower bounds, we take each pixel’s upper and lower quantiles.
Figure 2: Metric-guided bounds account for spatial correlations that affect downstream metrics. For maximum dose to the heart (D) with target coverage of 90%, we show contours for heart (red), right lung (blue), left lung (yellow), and body (green) overlaid on CT slices. Pixel-wise upper and lower bounds differ in pixel-wise intensity, while metric-guided bounds differ in the spatial distribution of pixel intensities. Pixel-wise upper bounds have larger heart volumes than lower bounds, while metric-guided bounds have similar heart volumes. Retrieval error $\varepsilon_{B}$ is the difference between estimated and actual bound divided by the interval length.
3.1 Validation
@@
We validate our method by computing coverage (Table 1), which is defined as the fraction of patients with ground truth metrics within the bounds. For metric-guided bounds, we use leave-one-out cross-validation on 20 patients and report the average coverage for metrics volume of ipsilateral lung that received 20Gy (Right Lung $V_{20}$), maxmimum dose to the heart (Heart $D_{0}$), and dose that 35% volume of the ipsilateral lung receives (Right Lung $D_{35}$). For conventional pixel-wise bounds, we compute the coverage of all patients. We use the finite sample correction $(1-\alpha)_{adj}=\frac{\lceil{(n_{p}+1)(1-\alpha)}\rceil}{n_{p}}$ [26, 2] for target coverage of $[(1-\alpha)_{adj}]$%. Our results show that metric-guided bounds give valid coverages for downstream tasks while conventional pixel-wise bounds do not.

### v2

@@ Abstract
Recent advancements in machine learning have led to the development of novel medical imaging systems and algorithms that address ill-posed problems. Assessing their trustworthiness and understanding how to deploy them safely at test time remains an important and open problem. In this work, we propose using conformal prediction to compute valid and distribution-free bounds on downstream metrics given reconstructions generated by one algorithm, and retrieve upper/lower bounds and inlier/outlier reconstructions according to the adjusted bounds. Our work offers 1) test time image reconstruction evaluation without ground truth, 2) downstream performance guarantees, 3) meaningful upper/lower bound reconstructions, and 4) meaningful statistical inliers/outlier reconstructions. We demonstrate our method on post-mastectomy radiotherapy planning using 3D breast CT reconstructions, and show 1) that metric-guided bounds have valid coverage for downstream metrics while conventional pixel-wise bounds do not and 2) anatomical differences of upper/lower bounds between metric-guided and pixel-wise methods. Our work paves way for more meaningful and trustworthy test-time evaluation of medical image reconstructions. Code available at https://github.com/matthewyccheung/conformal-metric.
Keywords Uncertainty  $\cdot$ Calibration  $\cdot$ Inverse Problems
@@ 1 Introduction
Image reconstruction is a crucial component of many clinical workflows. From computed tomography (CT) to magnetic resonance imaging (MRI), many medical imaging modalities are constructed by combining observed signals of patient anatomy into one “complete” image. For example, traditional CT reconstruction algorithms take thousands of 2D x-ray observations around a patient’s anatomy as input, and “backproject” these signals to produce 3D CT volumes [1]. These CT scans may in turn be used by radiotherapy treatment (RT) software to produce dosage plans for cancer treatment [2]. Image reconstruction accuracy and reliability can therefore have direct impacts on patient standard-of-care.
Often times, particularly in scenarios with limited imaging capabilities such as low-resource clinics [3, 4, 5], these reconstruction algorithms work with “sparse” observations, i.e., those that do not contain all information of the anatomy of interest. For example, sparse CT algorithms use limited ($<100$ instead of the standard $100$s) 2D x-ray observations to generate 3D CT scans [2, 6, 7]. In sparse image reconstruction, the observed information is not sufficient to recover the true image with complete certainty. Uncertainty originates from multiple sources in the reconstruction pipeline, including variations in image acquisition settings to quantum noise. To tackle uncertainty, classical reconstruction algorithms use human-defined priors, such as sparsity constraints [8], to choose one likely solution from the potentially infinite reconstructions consistent with the observations. While long successful, this approach does not describe the space of plausible reconstructions, and therefore, offers no insight into the resulting uncertainty in downstream clinical metrics which rely on those reconstructions, e.g., predicted radiation dose to organs in RT and organ volume.
Modern deep neural network (DNN) reconstruction algorithms, which learn complex priors from large datasets, offer the exciting alternative of returning multiple plausible reconstruction samples per input. Popular approaches to do so include deep generative modeling [9, 10], and randomized parameter initialization/dropout [11, 12]. A simple yet powerful way to use such samples is to compute confidence bounds on downstream clinical metrics. For example, in RT planning, confidence bounds can specify the range of doses to an organ , giving radiation physicists a better sense of the risks of the procedure.
While promising, there is a critical issue with using such bounds naively: if there are inherent biases of the reconstruction DNN, the resulting bounds will be misspecified. Indeed, DNNs are well-known to suffer from confidently incorrect predictions [13, 14], and in the context of generative tasks these are often termed “hallucinations” [15]. Examples of reconstruction hallucinations include systematically higher intensities for certain regions and certain organ geometry. These biases can degrade, and in the worst case, render confidence bounds on downstream metrics completely useless. Hence, there is a pressing need for methods to produce calibrated prediction bounds from DNN generative models, i.e., bounds that contain the ground truth metrics with certainty for new observations.
Recently, conformal prediction (CP) has gained acceptance as a powerful technique to calibrate uncertainty bounds of values (either categorical or continuous) predicted by a machine learning (ML) algorithm [16, 17, 18, 19, 20]. One approach is split conformalized quantile regression [21]. The main idea is to use a “calibration set” of data samples with ground truth target values to learn a compensation factor with which we may adjust bounds returned by the ML algorithm. Under the assumption of exchangeability, the resulting bounds will provably, on average, contain the ground truth target values $(1-\alpha)\%$ of the time for some set miscoverage rate $\alpha$ (e.g., often set to $0.05$ or $0.10$). However, while successful on a variety of traditional ML tasks such as image classification, the extension to image reconstruction problems is unnatural due the high-dimensional nature of the outputs.
In this work, we propose using CP to compute valid and distribution-free bounds on downstream metrics given images generated by an image reconstruction algorithm. Assuming a calibration set of reconstructions with true metric annotations (e.g., RT dose values to specific organs) as input, we use CP to adjust estimated quantiles of downstream metrics and return meaningful 1) upper/lower bound reconstructions and 2) statistical inliers/outliers.
We focus our empirical evaluations on post-mastectomy RT planning using 3D breast CT reconstructions, and show that 1) that metric-guided bounds have valid coverage for downstream metrics while conventional pixel-wise bounds do not and 2) anatomical differences of upper/lower bounds between metric-guided and pixel-wise methods.
Our work paves way for more meaningful and trustworthy evaluation of medical image reconstructions at test time.
1.1 Related Work
1.1.1 Conformal Prediction
Conformal prediction (CP) is a model-agnostic and distribution-free approach to attain valid prediction sets under the assumption of exchangeability [17, 18, 19, 20]. A common CP approach is split CP [16]. Based on a calibration dataset $\{(\hat{Y}^{1},Y^{1}),...,(\hat{Y}^{n},Y^{n})\}$, we can construct a prediction set $C(\hat{Y})$ for a fresh test point $Y^{n+1}$, such that:
$\mathbb{P}[Y^{n+1}\in C(\hat{Y}^{n+1})]\geq 1-\alpha$ (1)
where $\alpha$ is a user-specified miscoverage rate. This setup provides marginal coverage - on average, the prediction sets contain ground truth $(1-\alpha)\%$ of the time. For regression tasks, the prediction set is an interval [22]. Longer lengths indicate higher uncertainty. To derive the prediction intervals, the key idea is to use residuals from a calibration set as a conformity scores, $s$, to compute upper and lower bounds [17, 18, 19, 20]. For symmetric adjustments, the $(1-\alpha)$th quantile of the conformity scores is used to compute the prediction interval. For asymmetric adjustments, the conformity scores are computed separately for upper and lower adjustments. Many conformity scores exist, including absolute residual scores [22] and quantile-based scores [21]. Our work uses quantile-based conformity scores [21] to adjust estimated quantiles of downstream metrics from reconstruction samples to achieve valid prediction intervals.
1.1.2 Conformal Prediction for Image-to-Image Regression Problems
@@
Finding quantiles of high-dimensional data like images is a difficult problem because there are an infinite number of $(1-\alpha)$-quantiles and are only unique when a direction is specified [23, 24]. How do we pick such a direction? The conventional (pixel-wise) method is to pick the direction where all pixels are independent [25, 26, 27]. Recent work in CP provide upper and lower bounds based on pixel intensity [28, 29, 30] and principal components [31]. While pixel-wise prediction intervals are easy to interpret, they do not consider spatial correlations. While using principal components considers spatial correlations, it does not capture meaningful and practical uncertainty for downstream processes and is prohibitively costly to compute for large images. Furthermore, both methods provide upper and lower bounds not sampled from the learned manifold, yielding implausible images. An interesting approach is to calibrate the upper and lower bounds in the direction of semantic features [32]. However, this method requires training a generative model with disentangled latent spaces. Our work uses downstream metrics for calibration and retrieves image reconstruction bounds and prediction sets directly from the set of sampled reconstructions.
@@ 2 Method
Figure 1: Overview of our approach. Assume probabilistic measurement and reconstruction processes, $n_{p}$ patients for calibration, and 1 patient for testing. For test patient “$n_{p}+1$” with unknown ground truth reconstruction and metric, 1) acquire measurements, 2) attain a set of reconstructions, 3) extract downstream metrics, 4) adjust upper and lower bounds of metric based on a calibration procedure, and 5) retrieve reconstructions with metrics closest to the calibrated upper and lower bounds, contained within bounds (statistical inliers), and outside of bounds (statistical outliers).
We consider a 3-D reconstruction setting for a downstream application with a chosen downstream metric. The measurement and reconstruction algorithms are assumed to be probabilistic. We follow the split conformal prediction procedure [16] by using $n_{p}$ patients for the calibration dataset and 1 test patient $n_{p}+1$ with unknown ground truth volume and metric as the test dataset. For each patient $i$ in the calibration dataset, we reconstruct a set of volumes $\hat{V}^{i}=\{\hat{V}_{j}^{i}\}_{j=1}^{n_{r}}$ of size $n_{r}$. Each patient’s reconstructed volumes are used to attain a set of estimated metrics $\hat{Y}^{i}=\{\hat{Y}_{j}^{i}\}_{j=1}^{n_{r}}$. Each patient’s ground truth volume is used to attain a ground truth metric $Y^{i}$. For the test patient, we reconstruct a set of volumes $\hat{V}^{n_{p}+1}=\{\hat{V}_{j}^{n_{p}+1}\}_{j=1}^{n_{r}}$ and estimate metrics $\hat{Y}^{n_{p}+1}=\{\hat{Y}_{j}^{n_{p}+1}\}_{j=1}^{n_{r}}$. Assuming $(\hat{Y}^{i},Y^{i})$ for $i=1,...,n_{p}+1$ are exchangeable, we find the prediction interval satisfying the marginal coverage guarantee in Eq. 1 using a quantile-based conformity score $s^{i}=max[Q_{\alpha/2}(\hat{Y}^{i})-Y^{i},Y^{i}-Q_{1-\alpha/2}(\hat{Y}^{i})]$ [21]. The prediction interval is given as $C(\hat{Y}^{n_{p}+1})=[Q_{\alpha/2}(\hat{Y}^{n_{p}+1})-q,Q_{1-\alpha/2}(\hat{Y}^{n_{p}+1})+q]$ where $q$ is the $(1-\alpha)$th quantile of the conformity scores with a finite sample correction. Asymmetric bounds can be used as well, where the upper and lower bounds are adjusted separately. We retrieve the volumes 1) closest to the upper and lower bounds of the prediction intervals $[\hat{V}_{LB}^{n_{p}+1},\hat{V}_{UB}^{n_{p}+1}]$ based on the $L_{1}$ norm, 2) contained within the prediction intervals (inliers), and 3) outside the prediction intervals (outliers). We provide an overview in Fig. 1 and pseudo-code for symmetric bounds in Algorithm 1.
Algorithm 1 Metric-guided Retrieval (Symmetric Bounds)
$\triangleright$ Perform calibration to get upper and lower bound adjustment
@@
$q=Q_{\frac{\lceil(n_{p}+1)(1-\alpha)\rceil}{n_{p}}}(s)$
$\triangleright$ Compute prediction interval for patient in test dataset
@@
$\triangleright$ Retrieve upper and lower bound reconstructions
@@
$\triangleright$ Retrieve inliers and outlier reconstructions
$Inliers=\{\hat{V}_{j}^{n_{p}+1}\in\hat{V}^{n_{p}+1}|\hat{Y}_{j}^{n_{p}+1}\in[LB(\hat{Y}^{n_{p}+1}),UB(\hat{Y}^{n_{p}+1})]\}$
$Outliers=\{\hat{V}_{j}^{n_{p}+1}\in\hat{V}^{n_{p}+1}|\hat{Y}_{j}^{n_{p}+1}\notin[LB(\hat{Y}^{n_{p}+1}),UB(\hat{Y}^{n_{p}+1})]\}$
To verify the retrieved images are representative of the bounds at test time, we compute retrieval error defined as:
where $B$ denotes the calibrated bound and can be upper bound $UB$ or lower bound $LB$, and $\hat{Y}_{B}^{n_{p}+1}=\operatorname*{arg\,min}_{\hat{Y}_{j}^{n_{p}+1}}{|\hat{Y}_{j}^{n_{p}+1}-B^{n_{p}+1}|}$ are the estimated metrics closest to the calibrated bounds.
@@ 3 Experiments
Dataset: We use a de-identified CT dataset of 20 patients retrospectively treated with radiotherapy at The University of Texas MD Anderson Cancer Center. All CT images were of patients who had received surgical mastectomy to the right side of the body, and radiotherapy to the post-mastectomy chest wall and/or axillary lymph nodes. This research was conducted using an approved institutional review board protocol. Each ground truth CT is of size ($\times 512\times$ 512 × Number of slices). For each patient, we generate 10 digitally reconstructed radiographs (DRR) from the ground truth CT scan using the TIGRE toolbox [34]. The DRRs simulate image acquisition from a cone-beam geometry. We simulate physical randomness (beam angle variability and sensor noise) by generating DRRs with 3% noise and 50 random projections between 0 and 360 degrees. While simulating beam angle variability and sensor noise does not encompass all sources of randomness in practice, our method remains theoretically robust without requiring modifications to address additional factors of noise and bias. The number of projections was increased from 2 to 50 until organ boundaries were perceptually discernible in the reconstruction by the RPA. Because this work aims to showcase the feasibility of CP for image reconstruction, we assume that such a low-cost sparse CT device will be created in future work that gives acceptable reconstruction image quality. We use a self-supervised model, Neural Attenuation Fields (NAF), for reconstruction [35]. Each reconstruction is uncropped and contains the full scan. We use the default parameter setting in NAF [35] and introduce computational randomness through random initializations of NAF [36, 11]. Ultimately, we construct a dataset of 20 patients with 10 reconstructions each. To construct the conventional pixel-wise upper and lower bounds, we take each individual pixel’s upper and lower quantile intensities.
@@
Validation: We validate our method by assessing coverage, interval length, and number of outliers for our metric-guided method with symmetric and asymmetric adjustments, and the conventional pixel-wise method. Coverage is defined as the fraction of patients with ground truth metrics within the bounds. For metric-guided bounds, we use leave-one-out cross-validation on 20 patients and report metrics: volume of ipsilateral lung that received 20Gy (Right Lung $V_{20}$), maxmimum dose to the heart (Heart $D_{0}$), and dose that 35% volume of the ipsilateral lung receives (Right Lung $D_{35}$). For conventional pixel-wise bounds, we compute the coverage over all patients and process the pixel-wise upper and lower bounds in RPA to attain dose and structural statistics. We use the finite sample correction $(1-\alpha)_{adj}=\frac{\lceil{(n_{p}+1)(1-\alpha)}\rceil}{n_{p}}$ [21, 20] for target coverage of $[(1-\alpha)_{adj}]$%. If the lower bound is greater than the upper bound, the prediction interval size is zero and does not contain the ground truth.
4 Results
Our results are presented in Table. 1. We observe that while the average interval length of pixel-wise methods is smaller, pixel-wise methods suffer from significant undercoverage. In some cases like example Right Lung Volume, the average interval length is zero. This is because pixel-wise upper bounds produce smaller volumes than lower bounds. On the other hand, metric-guided bounds produce valid coverage. Our metric-guided method with symmetric adjustments tend to perform better than asymmetric adjustment (except for Right Lung Volume). We also show the probability of a reconstruction classified as outliers based on the bounds. While symmetric and asymmetric bounds perform similarly for Heart $D_{0}$, Heart Volume and Right Lung $V_{20}$, asymmetric bounds are more effective at identifying outliers for Right Lung Volume. We suspect these observations where asymmetric bounds are more effective than symmetric bounds are due to model misspecification (i.e. high estimation bias) since the reconstruction algorithm does not directly optimize for downstream metrics.
Method Metric Avg Test Coverage Avg Interval Length Avg P(outlier)
Metric-guided (sym) Heart $D_{0}$ (Gy) 0.90 4.39 0.09
Metric-guided (asym) Heart $D_{0}$ (Gy) 0.90 37.83 0.09
Pixel-wise Heart $D_{0}$ (Gy) 0.75 0.51 N/A
Metric-guided (sym) Heart Volume (cm3) 0.90 166.56 0.03
Metric-guided (asym) Heart Volume (cm3) 0.90 531.96 0.01
Pixel-wise Heart Volume (cm3) 0.50 81.46 N/A
Metric-guided (sym) Right Lung $V_{20}$ 0.90 0.07 0.18
Metric-guided (asym) Right Lung $V_{20}$ 0.90 0.09 0.17
Pixel-wise Right Lung $V_{20}$ 0.50 0.02 N/A
Metric-guided (sym) Right Lung $D_{35}$ (Gy) 0.90 9.13 0.14
Metric-guided (asym) Right Lung $D_{35}$ (Gy) 0.90 10.25 0.15
Pixel-wise Right Lung $D_{35}$ (Gy) 0.50 1.77 N/A
Metric-guided (sym) Right Lung Volume (cm3) 0.90 316.25 0.01
Metric-guided (asym) Right Lung Volume (cm3) 0.90 207.27 0.88
Pixel-wise Right Lung Volume (cm3) 0.00 0.00 N/A